# Soliton resolution for the energy-critical nonlinear heat equation in the radial case

### v2

@@ 2.3. Localized Energy Inequalities and Energy Trapping
$R^{(D-1)/2}v^{2}(R)$ $=-\int_{R}^{\infty}\partial_{r}(r^{(D-1)/2}v^{2}(r))\,\mathrm{d}r=-\int_{R}^{\infty}\left(\frac{D-1}{2r}v^{2}(r)+2v\partial_{r}v\right)r^{(D-1)/2}\,\mathrm{d}r$
$\leq 2\int_{R}^{\infty}|v||\partial_{r}v|r^{(D-1)/2}\,\mathrm{d}r$
$\leq 2\sqrt{\frac{1}{R^{D-3}}\int_{R}^{\infty}\frac{v^{2}}{r^{2}}r^{D-1}\,\mathrm{d}r\int_{R}^{\infty}|\partial_{r}v|^{2}r^{D-1}\,\mathrm{d}r}$
$\leq\frac{2}{R^{(D-3)/2}}\|v\|_{\mathcal{E}(R)}^{2}.$
Thus,
$|v(R)|\leq\frac{\sqrt{2}}{R^{(D-2)/2}}\|v\|_{\mathcal{E}(R)}.$
@@
$\begin{split}|v(r)|\leq\frac{\sqrt{2}}{r^{(D-2)/2}}\|v\|_{\mathcal{E}(r_{1},r_{2})}\end{split}$ (2.28)
@@ 3. Sequential bubbling for finite time blow-up solutions
See the works of Struwe [struwe] and Qing [qing] for analogous results for the harmonic map heat flow.
@@
where $\varepsilon_{0}>0$ is a constant coming from the $\varepsilon$–regularity established in Theorem 2.1 in [duShi]. Observe that when $u$ is radially symmetric then $\mathcal{S}=\{0\}$ since by radial Sobolev embedding (2.27) and $\sup_{t\in[0,T_{+})}\|u(t)\|_{\mathcal{E}}<\infty$ we have that $u\in L^{\infty}$ on the region $(r_{0},\infty)\times[0,T_{+})$ for any $r_{0}>0$. Since, the collection of maps $\{u(t)\}_{t\in[0,T_{+})}\subset\mathcal{E}$, it weakly converges (up to a subsequence) in $\mathcal{E}$ (or equivalently in $\dot{H}^{1}$) to $u^{*}\in\mathcal{E}.$ The uniqueness of $u^{*}$ follows from the fact that the weak limit $u^{*}\in\mathcal{E}$ is a weak stationary solution of (1.1) away from the origin with finite energy, and by standard removability of singularities argument (cf. [brezis1980removable]) one can show that it is a radial weak solution on $\mathbb{R}^{D}$. By elliptic regularity, one then concludes that it is a classical solution and is therefore equal to (1.6) up to symmetries.
Furthermore, since $u\in L^{\infty}$ on the region $0,T_{+})\times(r_{0},\infty)$ for any $r_{0}>0$, standard parabolic regularity theory implies $C^{\infty}$ convergence to a weak limit $u^{*}$ away from the origin, implying (3.1). Next to see (3.2), we argue as in the proof of Proposition 2.1 in [qing]. Denote,
@@
We will first show that $M=m.$ Note that $M$ and $m$ are independent of $r_{0}$ by (3.1). Thanks to (3.1) we can also assume that $r_{0}<\infty$. Take $\lambda_{j}=\frac{r_{0}}{j}$. Then observe that
@@
Then since $u(t)\rightarrow u^{*}$ smoothly outside the origin as $t\rightarrow T_{+}$, there exists a subsequence $\{t_{i_{j}}\}$ such that
@@
which combined with (LABEL:eqn:lambda_j-energy-expansion) and (3.6) implies
$\begin{split}\int_{0}^{\lambda_{j}}{\bf e}(u(t_{i_{j}}))r^{D-1}\mathrm{d}r&\geq M+\int_{0}^{r_{0}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-\int_{\lambda_{j}}^{r_{0}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-o_{j}(1)\\ &\geq M+\int_{0}^{\lambda_{j}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-o_{j}(1),\end{split}$ (3.11)
where $o_{j}(1)\to 0$ as $j\to\infty.$ Thus setting $t_{j}=t_{i_{j}}$ we get
@@
Up to subsequences we can arrange that $s_{j}\leq t_{j}\leq s_{j+1}\leq t_{j+1}$. Next, we derive a localized energy inequality for the localized nonlinear energy density ${\bf e}(u).$ Let $\phi\in C^{\infty}_{c}([0,\infty)).$ Then observe that
@@
Then using the above identity with $t_{1}=s_{j},t_{2}=t_{j}$ and $\phi$ such that $\phi\equiv 1$ on $0,\lambda_{j})$ and $\phi\equiv 0$ on $(\delta,\infty)$ for any $\delta\in(\lambda_{j},\infty)$ with $|\partial_{r}\phi|\lesssim(\delta-\lambda_{j})^{-1}$ we get
@@
where we use the fact that
@@
Consequently, sending $s_{j}\to T_{+}$ in (3.19) and using (3.12) we get
@@
Taking $\delta\rightarrow 0$, we can conclude that $m=M$. Thus, the limit $L=M$ exists in (3.2). Furthermore, $L$ is finite due to the fact that $\sup_{t\in[0,T_{+})}\|u(t)\|_{\mathcal{E}}<\infty.$
@@
Let $u_{0}\in\mathcal{E}$, and let $u(t)$ denote the solution to (1.1) with initial data $u_{0}$. Suppose that $T_{+}(u_{0})<\infty$. Then, there exist a unique function $u^{*}\in\mathcal{E}$, a unique integer $N\geq 1$, a sequence of times $t_{n}\to T_{+}$, signs $\vec{\iota}\in\{-1,1\}^{N}$, a sequence of scales $\vec{\lambda}_{n}\in(0,\infty)^{N}$, and an error $g_{n}$ defined by
@@
Using (2.15) with $|\partial_{r}\phi|\lesssim\frac{1}{r_{1}}$, for each $0<s<\tau<T_{+}$ we have
@@
$\tilde{\Theta}_{[r_{1},r_{2}]}(\tau)-\tilde{\Theta}^{*}_{[r_{1},r_{2}]}=\int_{r_{1}/2}^{2r_{2}}\left(\tilde{\mathbf{e}}(u(\tau))-\tilde{\mathbf{e}}(u^{*})\right)\phi(r)^{2}r^{D-1}\mathrm{d}r.$ (3.35)
@@
$\begin{split}&|\tilde{E}(u(s);\alpha(T_{+}-s)^{\frac{1}{2}},r_{0})-\tilde{E}(u^{*};0,r_{0})|\\ &\lesssim|\tilde{\Theta}_{[\alpha(T_{+}-s)^{\frac{1}{2}},r_{0}]}-\tilde{\Theta}^{*}_{[\alpha(T_{+}-s)^{\frac{1}{2}},r_{0}]}|+\tilde{E}(u(s);\alpha(T_{+}-s)^{\frac{1}{2}}/2,\alpha(T_{+}-s)^{\frac{1}{2}})\\ &\quad+\tilde{E}(u^{*};0,\alpha(T_{+}-s)^{\frac{1}{2}})+|\int_{r_{0}}^{2r_{0}}(\tilde{{\bf e}}(u(s))-\tilde{{\bf e}}(u^{*}))\phi^{2}r^{D-1}\mathrm{d}r|\end{split}$ (3.40)
@@
We now aim to show that $L=NE(W)$ and that $N\geq 1.$ If $N=0$, then from the decomposition (3.30) we deduce that $\|u(t_{n})\|_{\mathcal{E}(r\leq\alpha(T_{+}-t_{n})^{\frac{1}{2}})}\to 0$ as $n\to\infty$ which implies that $0\not\in\mathcal{S}$ and therefore by $\varepsilon$-regularity $u(t)$ is in $L^{\infty}$ for $r\in[0,\infty)$ up until $t=T_{+}$ contradicting the definition of the maximal blow up time (cf. Section 2.6 in [collotdescription]). Thus $N\geq 1.$ To see that $L=NE(W)$ we observe that along the sequence $\{t_{n}\}$ in (3.30) we have the following energy identity:
@@ 4. Sequential bubbling for global solutions
Let $u_{0}\in\mathcal{E}$ and let $u(t)$ denote the solution to (1.1) with initial data $u_{0}$. Suppose that $T_{+}(u_{0})=\infty$. Then there exist $T_{0}\in(0,\infty)$, a unique integer $N\geq 0$, a sequence of times $t_{n}\to\infty$, signs $\vec{\iota}\in\{-1,1\}^{N}$, a sequence of scales $\vec{\lambda}_{n}\in(0,\infty)^{N}$, and a sequence of $g_{n}$ defined by
@@
$\phi(t,r)=\phi_{T}(r)=1-\chi(4r/\alpha\sqrt{T}){\ \ \text{for}\ \ }t\in[T_{0},T]$
where $\chi(r)$ is a smooth function on $(0,\infty)$ such that $\chi(r)=1$ for $r\leq 1$, $\chi(r)=0$ if $r\geq 4$, and $\left\lvert{\chi^{\prime}(r)}\right\rvert\leq 1$ for all $r\in(0,\infty)$. Since $\frac{\mathrm{d}}{\mathrm{d}t}\phi(t,r)=0$ for $t\in[T_{0},T]$ it follows from (2.25) that
@@
where we used $|\partial_{r}\phi|^{2}\lesssim T^{-1}$ and the radial Sobolev embedding 2.27 to control the integral involving $|u|^{2p}$. The constant in the above inequality depends on $\sup_{t\in[0,\infty)}\|u(t)\|_{\mathcal{E}}$ and $\alpha$. So in particular we can make this small by choosing $T_{0}$ large enough such that $\tilde{C}\left(\int_{T_{0}}^{\infty}\int_{0}^{\infty}(\partial_{t}u)^{2}\right)^{1/2}\leq\frac{\varepsilon}{2}.$ Using the above together with (4.7) and (4.8), we find that the Dirichlet energy can be made arbitrarily small
@@ 5. Decomposition of the solution and collision intervals
For the remainder of the paper, we fix a solution $u(t)\in\mathcal{E}$ of (1.1), defined on the time interval $I_{*}=[0,T_{+})$ where $T_{+}<\infty$ or $T_{+}=\infty$. By Propositions 3.3 and 4.1 there exist a unique integer $N\geq 0$, a unique function $u^{*}\in\mathcal{E}$ and a sequence of times $t_{n}\to T_{+}$ so that $u(t_{n})-u^{*}$ approaches an $N$-bubble as $n\to\infty$. When $T_{+}=\infty$, we simply set $u^{*}=0$ and when $T_{+}<\infty$ then $N\geq 1$. We define the localized distance to an $N$-bubble.
@@
Note that it suffices to prove Theorem 1.1 in the case when $N\geq 1$. Indeed when $N=0$, $T_{+}=\infty$ by Proposition (3.3) and $\lim_{n\to\infty}\|u(t_{n})\|_{\mathcal{E}}=0$ by Proposition (4.1), so the small data Cauchy theory from Lemma 2.1 implies $\lim_{t\to\infty}\|u(t)\|_{\mathcal{E}}=0$. Furthermore, as a direct consequence of (3.27) and (4.4), defining $\rho_{N}(t)=\rho(t)$ gives us the following lemma.
@@ 5.1. Collision intervals
To prove Theorem 1.1 we analyze collision intervals. We recall their definition from [lawrie-harmonic-map].
@@
The proof is similar to the proof of Lemma 5.6 in [lawrie-wave], but we recall it here for the reader’s convenience.
First, we will show that $K$ is well defined with $K\leq N$. It suffices to show that there exist $\eta>0$, a decreasing sequence $\varepsilon_{n}\to 0$, and sequences $(a_{n}),(b_{n})$ such that $a_{n},b_{n}]\in\mathcal{C}(\varepsilon_{n},\eta)$ for all $n\in\mathbb{N}.$ Since (5.3) is false, there exist $\eta>0$ and a monotone sequence $b_{n}\to T_{+}$ such that for all $n$,
@@ 5.2. Decomposition of the solution
for all $t\in\left[a_{n},b_{n}\right]$ and all $n\in\mathbb{N}$. Furthermore, for any sequence $s_{n}\in\left[a_{n},b_{n}\right]$ and any sequence $R_{n}$ such that $\nu\left(s_{n}\right)\leq R_{n}\ll\lambda_{K+1}\left(s_{n}\right)$ if $K<N$ and $\nu\left(s_{n}\right)\leq R_{n}\leq\alpha\left(s_{n}\right)$ if $K=N$,
$\begin{split}\lim_{n\rightarrow\infty}{E}\left(u\left(s_{n}\right);R_{n},\infty\right)=(N-K){E}(Q)+{E}\left(u^{*}\right)\end{split}$ (5.28)
@@
$\begin{split}\lim_{n\rightarrow\infty}\left(\|g^{\mathrm{ext}}_{n}\|_{\mathcal{E}(r\geq R_{n})}+\sum_{j=K+1}^{N}\left(\frac{\lambda_{j}(s_{n})}{\lambda_{j+1}(s_{n})}\right)^{\frac{D-2}{4}}\right)=0.\end{split}$ (5.29)
@@
We closely following the proof of Lemma 5.9 in [lawrie-harmonic-map] and restrict ourselves to the case when $T_{+}<\infty$ since the case when $T_{+}=\infty$ is similar and less cumbersome due to the fact that $u^{*}=0.$
@@
Next, we define the scales $\alpha(t)$ and $\nu(t)$ as in (5.19) and (5.16) respectively. To this end, using Definition 5.1, for all $n\in\mathbb{N}$, there exist scales scales $\rho_{K}(t)\ll\mu_{K+1}(t)\ll\cdots\ll\mu_{N}(t)\ll\left(T_{+}-t\right)^{\frac{1}{2}}$ and signs $\vec{\sigma}(t)\in\{-1,1\}^{N-K}$ for $t\in\left[a_{n},b_{n}\right]$, such that the error function $\tilde{g}_{\rho_{K}}(t)$ defined on the region $r\in\left(\rho_{K}(t),\infty\right)$ by the identity
@@
where the first inequality follows by integration by parts and Cauchy-Schwarz, the second inequality follows from a computation analogous to Lemma 2.21 in [lawrie-wave], the third inequality follows from a Taylor expansion, and the fourth inequality follows from (5.25). Thus, collecting all the above estimates, we get

### v3

@@ 2.3. Localized Energy Inequalities and Energy Trapping
$R^{D-2}v^{2}(R)$ $=\int_{R}^{\infty}\left(-\frac{D-2}{r}v^{2}(r)-2v\partial_{r}v\right)r^{D-2}\mathrm{~d}r$
$\leq\int_{R}^{\infty}2|v|\left|\partial_{r}v\right|r^{D-2}\mathrm{~d}r\leq\int_{R}^{\infty}\left(\frac{|v|^{2}}{r^{2}}+\left|\partial_{r}v\right|^{2}\right)r^{D-1}\mathrm{~d}r=\|v\|_{\mathcal{E}(R)}^{2},$
which on re-arranging implies (2.27). ∎
@@
$\begin{split}|v(r)|\leq\frac{1}{r^{(D-2)/2}}\|v\|_{\mathcal{E}(r_{1},r_{2})}\end{split}$ (2.28)
@@ 3. Sequential bubbling for finite time blow-up solutions
See the works of Struwe [STR85] and Qing [QIN95] for analogous results for the harmonic map heat flow.
@@
[…]
Furthermore, since $u\in L^{\infty}$ on the region $0,T_{+})\times(r_{0},\infty)$ for any $r_{0}>0$, standard parabolic regularity theory implies $\dot{H}^{1}\cap C^{0}$ convergence to a weak limit $u^{*}$ away from the origin, implying (3.1). Next to see (3.2), we argue as in the proof of Proposition 2.1 in [QIN95]. Denote,
@@
It suffices to show $M=m.$ Note that $M$ and $m$ are independent of $r_{0}$ by (3.1). Thanks to (3.1) we can also assume that $r_{0}<\infty$. Take $\lambda_{j}=\frac{r_{0}}{j}$. Then observe that
@@
Then since $u(t)\rightarrow u^{*}$ smoothly outside the origin as $t\rightarrow T_{+}$, there exist subsequences still denoted by $\{t_{i_{j}}\}$ and $\{s_{i^{\prime}_{j}}\}$ such that
@@
and
$\begin{split}\int_{\lambda_{j}}^{r_{0}}|{\bf e}(u(s_{i^{\prime}_{j}}))-{\bf e}(u^{*})|r^{D-1}\mathrm{d}r\to 0,\quad\text{ as }j\to\infty.\end{split}$ (3.11)
Denote $t_{j}=t_{i_{j}}$ and $s_{j}=s_{i^{\prime}_{j}}$. Up to subsequences we can arrange that $s_{j}\leq t_{j}\leq s_{j+1}\leq t_{j+1}$. Then (3.10), (LABEL:eqn:lambda_j-energy-expansion) and (3.6) implies
$\begin{split}\int_{0}^{\lambda_{j}}{\bf e}(u(t_{j}))r^{D-1}\mathrm{d}r&\geq M+\int_{0}^{r_{0}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-\int_{\lambda_{j}}^{r_{0}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-o_{j}(1)\\ &\geq M+\int_{0}^{\lambda_{j}}{\bf e}(u^{*})r^{D-1}\mathrm{d}r-o_{j}(1),\end{split}$ (3.12)
where $o_{j}(1)\to 0$ as $j\to\infty.$ Thus,
@@
Next, we derive a localized energy inequality for the localized nonlinear energy density ${\bf e}(u).$ Let $\phi\in C^{\infty}_{c}([0,\infty)).$ Then observe that
@@
Then using the above identity with $t_{1}=s_{j},t_{2}=t_{j}$ and $\phi$ such that $\phi\equiv 1$ on $0,\lambda_{j})$ and $\phi\equiv 0$ on $(\delta,\infty)$ for any $\delta\in(\lambda_{j},r_{0})$ with $|\partial_{r}\phi|\lesssim(\delta-\lambda_{j})^{-1}$ we get
@@
where we use (3.11) and a similar argument as above to get that
@@
Consequently, sending $s_{j}\to T_{+}$ and recalling in (3.20) and using (3.13) we get
@@
Taking $\delta\rightarrow 0$, and recalling (3.8), we can conclude that $m=M$. Thus, the limit $L=M$ exists in (3.2). Furthermore, $L$ is finite due to the fact that $\sup_{t\in[0,T_{+})}\|u(t)\|_{\mathcal{E}}<\infty.$
@@
Let $u_{0}\in\mathcal{E}$, and let $u(t)$ denote the solution to (1.1) with initial data $u_{0}$ with bounded $\mathcal{E}$-norm. Suppose that $T_{+}(u_{0})<\infty$. Then, there exist a unique function $u^{*}\in\mathcal{E}$, a unique integer $N\geq 1$, a sequence of times $t_{n}\to T_{+}$, signs $\vec{\iota}\in\{-1,1\}^{N}$, a sequence of scales $\vec{\lambda}_{n}\in(0,\infty)^{N}$, and an error $g_{n}$ defined by
@@
Using (2.25), $|\partial_{r}\phi|\lesssim\frac{1}{r_{1}}$, and (2.27) for each $0<s<\tau<T_{+}$ we have
@@
$\tilde{\Theta}_{[r_{1},r_{0}]}(\tau)-\tilde{\Theta}^{*}_{[r_{1},r_{0}]}=\int_{r_{1}/2}^{2r_{0}}\left(\tilde{\mathbf{e}}(u(\tau))-\tilde{\mathbf{e}}(u^{*})\right)\phi(r)^{2}r^{D-1}\mathrm{d}r.$ (3.36)
@@
$\begin{split}&|\tilde{E}(u(s);\alpha(T_{+}-s)^{\frac{1}{2}},r_{0})-\tilde{E}(u^{*};0,r_{0})|\\ &\lesssim|\tilde{\Theta}_{[\alpha(T_{+}-s)^{\frac{1}{2}},r_{0}]}(s)-\tilde{\Theta}^{*}_{[\alpha(T_{+}-s)^{\frac{1}{2}},r_{0}]}|+\tilde{E}(u(s);\alpha(T_{+}-s)^{\frac{1}{2}}/2,\alpha(T_{+}-s)^{\frac{1}{2}})\\ &\quad+\tilde{E}(u^{*};0,\alpha(T_{+}-s)^{\frac{1}{2}})+|\int_{r_{0}}^{2r_{0}}(\tilde{{\bf e}}(u(s))-\tilde{{\bf e}}(u^{*}))\phi^{2}r^{D-1}\mathrm{d}r|\end{split}$ (3.41)
@@
We now aim to show that $L=NE(W)$ and that $N\geq 1.$ If $N=0$, then from the decomposition (3.31) we deduce that $\|u(t_{n})\|_{\mathcal{E}(r\leq\alpha(T_{+}-t_{n})^{\frac{1}{2}})}\to 0$ as $n\to\infty$ which along with (3.27), implies that $0\not\in\mathcal{S}$ and therefore by $\varepsilon$-regularity $u(t)$ is in $L^{\infty}$ for $r\in[0,\infty)$ up until $t=T_{+}$ contradicting the definition of the maximal blow up time (cf. Section 2.6 in [5]). Thus $N\geq 1.$ To see that $L=NE(W)$ we observe that along the sequence $\{t_{n}\}$ in (3.31) we have the following energy identity:
@@ 4. Sequential bubbling for global solutions
Let $u_{0}\in\mathcal{E}$ and let $u(t)$ denote the solution to (1.1) with initial data $u_{0}$ and bounded $\mathcal{E}$-norm. Suppose that $T_{+}(u_{0})=\infty$. Then there exist $T_{0}\in(0,\infty)$, a unique integer $N\geq 0$, a sequence of times $t_{n}\to\infty$, signs $\vec{\iota}\in\{-1,1\}^{N}$, a sequence of scales $\vec{\lambda}_{n}\in(0,\infty)^{N}$, and a sequence of $g_{n}$ defined by
@@
$\phi(t,r)=\phi_{T}(r)=1-\chi(4r/\alpha\sqrt{T}){\ \ \text{for}\ \ }t\in[T_{0},T].$
Since $\frac{\mathrm{d}}{\mathrm{d}t}\phi(t,r)=0$ for $t\in[T_{0},T]$ it follows from (2.25) that
@@
where we used $|\partial_{r}\phi|^{2}\lesssim T^{-1}$ and the radial Sobolev embedding (2.27) to control the integral involving $|u|^{2p}$. The constant in the above inequality depends on $\sup_{t\in[0,\infty)}\|u(t)\|_{\mathcal{E}}$ and $\alpha$. So in particular we can make this small by choosing $T_{0}$ large enough such that $\tilde{C}\left(\int_{T_{0}}^{\infty}\int_{0}^{\infty}(\partial_{t}u)^{2}\right)^{1/2}\leq\frac{\varepsilon}{2}.$ Using the above together with (4.7) and (4.8), we find that the Dirichlet energy can be made arbitrarily small
@@ 5. Decomposition of the solution and collision intervals
For the remainder of the paper, we fix a solution $u(t)$ of (1.1), with bounded energy (in $\mathcal{E}$-norm) defined on the time interval $I_{*}=[0,T_{+})$ where $T_{+}<\infty$ or $T_{+}=\infty$. By Propositions 3.3 and 4.1 there exist a unique integer $N\geq 0$, a unique function $u^{*}\in\mathcal{E}$ and a sequence of times $t_{n}\to T_{+}$ so that $u(t_{n})-u^{*}$ approaches an $N$-bubble as $n\to\infty$. When $T_{+}=\infty$, we simply set $u^{*}=0$ and when $T_{+}<\infty$ then $N\geq 1$. We define the localized distance to an $N$-bubble.
@@
Note that it suffices to prove Theorem 1.1 in the case when $N\geq 1$. Indeed when $N=0$, $T_{+}=\infty$ by Proposition 3.3 and $\lim_{n\to\infty}\|u(t_{n})\|_{\mathcal{E}}=0$ by Proposition 4.1, so the small data Cauchy theory from Lemma 2.1 implies $\lim_{t\to\infty}\|u(t)\|_{\mathcal{E}}=0$. Furthermore, as a direct consequence of (3.28) and (4.4), defining $\rho_{N}(t)=\rho(t)$ gives us the following lemma.
@@ 5.1. Collision intervals
To prove Theorem 1.1 we analyze collision intervals. We recall their definition from [JL23a].
@@
The proof is similar to the proof of Lemma 5.6 in [JL23b], but we recall it here for the reader’s convenience.
First, we will show that $K$ is well defined with $K\leq N$. It suffices to show that there exist $\eta>0$, a decreasing sequence $\varepsilon_{n}\to 0$, and sequences $(a_{n}),(b_{n})$ such that $a_{n},b_{n}]\in\mathcal{C}_{N}(\varepsilon_{n},\eta)$ for all $n\in\mathbb{N}.$ Since (5.3) is false, there exist $\eta>0$ and a monotone sequence $b_{n}\to T_{+}$ such that for all $n$,
@@ 5.2. Decomposition of the solution
for all $t\in\left[a_{n},b_{n}\right]$ and all $n\in\mathbb{N}$. Furthermore, for any sequence $s_{n}\in\left[a_{n},b_{n}\right]$
$\begin{split}\lim_{n\rightarrow\infty}{E}\left(u\left(s_{n}\right);\nu(s_{n}),\infty\right)=(N-K){E}(Q)+{E}\left(u^{*}\right)\end{split}$ (5.28)
@@
$\begin{split}\lim_{n\rightarrow\infty}\left(\|g^{\mathrm{ext}}_{n}\|_{\mathcal{E}(r\geq\nu(s_{n}))}+\sum_{j=K+1}^{N}\left(\frac{\lambda_{j}(s_{n})}{\lambda_{j+1}(s_{n})}\right)^{\frac{D-2}{4}}\right)=0.\end{split}$ (5.29)
@@
We closely follow the proof of Lemma 5.9 in [JL23a] and restrict ourselves to the case when $T_{+}<\infty$ since the case when $T_{+}=\infty$ is similar and less cumbersome due to the fact that $u^{*}=0.$
@@
Next, we define the scales $\alpha(t)$ and $\nu(t)$ as in (5.19) and (5.16) respectively. To this end, using Definition 5.1, for all $n\in\mathbb{N}$, there exist scales $\rho_{K}(t)\ll\mu_{K+1}(t)\ll\cdots\ll\mu_{N}(t)\ll\left(T_{+}-t\right)^{\frac{1}{2}}$ and signs $\vec{\sigma}(t)\in\{-1,1\}^{N-K}$ for $t\in\left[a_{n},b_{n}\right]$, such that the error function $\tilde{g}_{\rho_{K}}(t)$ defined on the region $r\in\left(\rho_{K}(t),\infty\right)$ by the identity
@@
where the first inequality follows by integration by parts and Cauchy-Schwarz, the second inequality follows from a computation analogous to Lemma 2.21 in [JL23b], the third inequality follows from a Taylor expansion, and the fourth inequality follows from (5.25). Thus, collecting all the above estimates, we get